OpenStudy (meilendurcer):

What is the slope of the line that contains (3, 4) and (–5, 2)? A. 14/56 B. -4 C. 4 D. -14/56

OpenStudy (anonymous):

Slope equation: \[m = \frac{ y _{2}- y _{1} }{ x _{2}- x1}\] Your points: (3, 4) and (–5, 2) Where Y2= 2 and Y1= 4 And X2 = -5 and X1=3 Now plug it in and solve

OpenStudy (meilendurcer):

@zaynahf -4?

OpenStudy (anonymous):

How did you get that?

OpenStudy (meilendurcer):

i guessed.since all i got was -2

OpenStudy (anonymous):

How did you get from -2 to -4..

OpenStudy (meilendurcer):

-2 - 2 =-4

OpenStudy (anonymous):

This is what you're supposed to be doing: \[\frac{ 2-4 }{ -5 - 3 }\]

OpenStudy (anonymous):

Subtract 2-4= -2 and -5-3=-8 -2/-8 is your slope

OpenStudy (anonymous):

Simplify it which is 1/4 =m (slope)